A body is thrown up with the velocity of 15 metre per second then the maximum height attained by the body is

Grade:11

2 Answers

Arun
6 years ago
Dear student
 

Maximum height of a vertically projected body is given by h=u^2/2g . for derivation see the other answer.

So the answer will be h= 15^2/2g

= 225/20 (taking g = 10 m/s^2)

= 11.25 m

 

Regards

Arun (askIITians forum expert)

RJ 123456
6 years ago
Maximum height is given by, h=u2/2g, where is maximum height, is initial velocity and is acceleration due to gravity.
So, Maximum Height, h=152/2g
                                         =225/2×10.  (g=10m/s2)
                                         =11.25 m